OpenStudy (anonymous):

2(x + 6) − (x + 1) = 16

OpenStudy (anonymous):

2x + 12 - x - 1 = 16 x + 11 = 16 Isn't it simple now?

OpenStudy (anonymous):

subtract 11

OpenStudy (anonymous):

an x = 5

OpenStudy (anonymous):

x is 5

OpenStudy (anonymous):

You're right!

OpenStudy (anonymous):

open the brackets 2x +12 - x - 1 = 16 x = 5